OpenStudy (anonymous):

Medal /Fan the volumes of two similar solids are 729m^3 and 125m^3. The surface area of the larger solid is 324m^2. What is the surface area of the smaller solid?

OpenStudy (anonymous):

@mathslover could you help me ?

mathslover (mathslover):

See, 729 = 9^3 and 125 = 5^3 Right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

So, here I will consider the volume as the function of sides of the solid. \(\cfrac{V_1}{V_2} = \cfrac{9^3}{5^3} = \cfrac{{L_1} ^3}{{L_2}^3}\) So, the sides are in the ratio of 9:5 Surface area will be in the ratio of : \(\cfrac{{L_1}^2}{{L_2}^2} \) Can you move on from here?

OpenStudy (anonymous):

yes, thank you

mathslover (mathslover):

Great, welcome to OpenStudy and good luck!